# Capturing Turbulence with Numerical Dissipation: a Simple Dynamical Model for Unresolved Turbulence in Hydrodynamic Simulations

### v2

@@ 5 Discussion
Another potential reason why the SLES and explicit LES models produce different results in the subsonic regime of the disk outskirts is numerical. In this regime, $e_{\rm turb}$ becomes relatively small compared to other energy terms in Equation 2 making the estimate numerically unreliable. In fact, the positivity of the implicit source term in that equation is not generally guaranteed. To guarantee this positivity, the hydrodynamic method must be entropy-stable, i.e., the local entropy must not decrease, or there must be no negative dissipation (e.g., tadmor03; schmidtmann17). In our turbulent box tests, we find that such “negative dissipation” regions occur in the low-density gas next to strong density gradients but they constitute a negligible fraction of the total volume and energy budget. Most of the dissipation is positive and it occurs in high-density converging flows, leading to a strong positive correlation of $e_{\rm turb}$ with gas density (recall Figure 2). These effects, however, can become important for modeling subgrid turbulence in subsonic, weakly compressible regimes, which we leave to a separate study.
@@
Apart from shearing flows, another example where the dissipated energy can bypass a turbulent cascade and be thermalized directly is strong shocks. To this end, one can use a shock-finding algorithm and make the fraction of thermalized energy dependent on the presence and local properties of the shock. In our current implementation of the SLES model, this is not taken into account, and strong shocks convert kinetic energy into $e_{\rm turb}$ instead. Note, however, that this does not affect the shock solution because the total post-shock pressure remains identically the same, independent of the partitioning between $e_{\rm turb}$ and $e_{\rm th}$ as both are described by the ideal-gas equation of state with the same adiabatic index of $5/3$. In addition, our tests in the decaying supersonic turbulence setup show that our more simple model works reasonably well (Section 3).
@@ 6 Summary
We have introduced a dynamic model for unresolved turbulence in hydrodynamic simulations that explicitly tracks subgrid turbulent energy by sourcing it implicitly via numerically dissipated kinetic energy in each cell. This model is physically motivated by a generic property of turbulent flows that they dissipate kinetic energy at a rate set by the energy cascade rate from large scales, which is independent of fluid viscosity regardless of its nature, be it physical or numerical. For this reason, the local numerical dissipation rate can be used as a reasonable approximation for the local rate of energy cascade from resolved to subgrid scales.
@@
In its simplest form presented here, this model involves only one tunable parameter that describes subgrid turbulence decay into heat. We calibrate this parameter against high-resolution simulations of decaying supersonic turbulence (Section 3.2) and then test its predictions in a turbulent box (Section 3.3) and isolated galaxy disk simulations (Section 4). Despite its simplicity, the model can capture many non-trivial properties of these turbulent flows. These results are summarized as follows:
@@
    Small-scale turbulence exhibits two limiting regimes of density-dependence. In weak turbulence, cell-crossing time is long, and therefore turbulence decays slowly and behaves near-adiabatically ($e_{\rm turb}\propto\rho^{5/3}$, or $\sigma_{\rm turb}\propto\rho^{1/3}$). In strong turbulence, cell-crossing time is short, and turbulence behaves near-isothermally ($e_{\rm turb}\propto\rho$, or $\sigma_{\rm turb}={\rm const}$). The $\sigma_{\rm turb}$$\rho$ relation slope in our decaying supersonic turbulence test exhibits intermediate values, indicating that the small-scale turbulence is, on average, in an intermediate regime (Figure 5).
@@
Our results indicate that the new SLES model can produce realistic results—at least in the context of modeling supersonic star-forming ISM—at a fraction of the cost of explicit LES. The model is significantly more straightforward to implement in existing codes used to carry out implicit LES as it largely relies on the already implemented infrastructure. This warrants further investigation of the model performance in other regimes (such as shearing and subsonic flows pointed out above) and different hydrodynamic codes.
@@ Acknowledgements
I am deeply grateful to Christoph Federrath, Troels Haugbølle, Lars Hernquist, Ralf Klessen, Andrey Kravtsov, Alexei Kritsuk, Jonathan Stern, Romain Teyssier, and the participants of the Ringberg workshop “Advancements in Computational Galaxy Formation” for insightful discussions. I also thank Alexei Kritsuk for providing access to the initial conditions of an isothermal turbulence test used in the code comparison project of kritsuk11. The analyses and simulations presented in this paper were performed on the NASA Pleiades cluster supported by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center and on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University. Support for V.S. was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51445.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555, and by Harvard University through the Institute for Theory and Computation Fellowship. The analyses presented in this paper were greatly aided by the following free software packages: NumPy (numpy_ndarray), SciPy (scipy), Matplotlib (matplotlib), and yt (yt). We have also used the Astrophysics Data Service (ADS) and arXiv preprint repository extensively during this project and the writing of the paper.
@@ Appendix A Dependence on resolution
The turbulent cascade in simulations without explicit viscosity truncates at the numerical dissipation scale close to resolution. As Figure 7 shows, in our high-resolution ($1024^{3}$) DNS that we use to calibrate and test our SLES model this truncation starts around $k/k_{\rm min}=L_{\rm box}/L\sim 100$, which corresponds to $\sim$10 cells in agreement with kritsuk11 results. Only the estimates of small-scale turbulence, $e_{\rm turb}$, on sufficiently larger scales should be used for comparisons with SLES, as $e_{\rm turb}$ on smaller scales is not numerically converged.
In Figure 12, we investigate the appropriate coarse-graining scales that can be used for our purposes. Different line styles show DNS with different resolutions, while the line colors show the average $E_{\rm turb}\equiv\langle e_{\rm turb}\rangle_{\rm box}$ coarse-grained using different coarse grids: $16^{3}$, $32^{3}$, $64^{3}$, $128^{3}$ (see Section 3.1). A finer coarse grid corresponds to the estimate of $e_{\rm turb}$ over a smaller number of mesh cells; e.g., for our fiducial $1024^{3}$ DNS, these grids correspond to 64, 32, 16, and 8 cells per side.
The comparison of DNS with different resolutions and the spacing between the results for different coarse grids shows, the results of our DNS run are converged when considered on down to $64^{3}$ coarse grids. In Section 3, however, we also show the results for the $128^{3}$, which however should be interpreted with caution as they probe $e_{\rm turb}$ on scales close to the numerical dissipation scale.

### v3

@@ 5 Discussion
Another potential reason why the SLES and explicit LES models produce different results in the subsonic regime of the disk outskirts is numerical. In this regime, $e_{\rm turb}$ becomes relatively small compared to other energy terms in Equation (2), making the estimate numerically unreliable. In fact, the positivity of the implicit source term in that equation is not generally guaranteed. To guarantee this positivity, the hydrodynamic method must be entropy-stable, i.e., the local entropy must not decrease, or there must be no negative dissipation (e.g., tadmor03; schmidtmann17). In our turbulent box tests, we find that such “negative dissipation” regions occur in the low-density gas next to strong density gradients, but they constitute a negligible fraction of the total volume and energy budget. Most of the dissipation is positive, and it occurs in high-density converging flows, leading to a strong positive correlation of $e_{\rm turb}$ with gas density (recall Figure 2). These effects, however, can become important for modeling subgrid turbulence in subsonic, weakly compressible regimes, which we leave to a separate study.
@@
Apart from shearing flows, another example where the dissipated energy can bypass a turbulent cascade and be thermalized directly is strong shocks. To this end, one can use a shock-finding algorithm and make the fraction of thermalized energy dependent on the presence and local properties of the shock. In our current implementation of the SLES model, this is not taken into account, and strong shocks convert kinetic energy into $e_{\rm turb}$ instead. Note, however, that this does not affect the shock solution because the total postshock pressure remains identically the same, independently of the partitioning between $e_{\rm turb}$ and $e_{\rm th}$ as both are described by the ideal gas equation of state with the same adiabatic index of $5/3$. In addition, our tests in the decaying supersonic turbulence setup show that our simpler model works reasonably well (Section 3).
@@ 6 Summary
We have introduced a dynamic model for unresolved turbulence in hydrodynamic simulations that explicitly tracks subgrid turbulent energy by sourcing it implicitly via numerically dissipated kinetic energy in each cell. This model is physically motivated by a generic property of turbulent flows that they dissipate kinetic energy at a rate set by the energy cascade rate from large scales, which is independent of fluid viscosity, regardless of its nature, be it physical or numerical. For this reason, the local numerical dissipation rate can be used as a reasonable approximation for the local rate of energy cascade from resolved to subgrid scales.
@@
In its simplest form, presented here, this model involves only one tunable parameter that describes subgrid turbulence decay into heat. We calibrate this parameter against high-resolution simulations of decaying supersonic turbulence (Section 3.2) and then test its predictions in a turbulent box (Section 3.3) and isolated galaxy disk simulations (Section 4). Despite its simplicity, the model can capture many nontrivial properties of these turbulent flows. These results are summarized as follows:
@@
    Small-scale turbulence exhibits two limiting regimes of density dependence. In weak turbulence, cell-crossing time is long, and therefore turbulence decays slowly and behaves near-adiabatically ($e_{\rm turb}\propto\rho^{5/3}$, or $\sigma_{\rm turb}\propto\rho^{1/3}$). In strong turbulence, cell-crossing time is short, and turbulence behaves near-isothermally ($e_{\rm turb}\propto\rho$, or $\sigma_{\rm turb}={\rm const}$). The $\sigma_{\rm turb}$$\rho$ relation slope in our decaying supersonic turbulence test exhibits intermediate values, indicating that the small-scale turbulence is, on average, in an intermediate regime (Figure 5).
@@
Our results indicate that the new SLES model can produce realistic results—at least in the context of modeling supersonic star-forming ISM—at a fraction of the cost of explicit LES. The model is significantly more straightforward to implement in existing codes used to carry out ILES as it largely relies on the already implemented infrastructure. This warrants further investigation of the model performance in other regimes (such as shearing and subsonic flows pointed out above) and different hydrodynamic codes.
@@ Acknowledgements
I would like to thank the anonymous referee for constructive comments, which helped improve the manuscript. I am also grateful to Christoph Federrath, Troels Haugbølle, Lars Hernquist, Ralf Klessen, Andrey Kravtsov, Alexei Kritsuk, Jonathan Stern, Romain Teyssier, and the participants of the Ringberg workshop “Advancements in Computational Galaxy Formation” for insightful discussions. I also thank Alexei Kritsuk for providing access to the initial conditions of an isothermal turbulence test used in the code comparison project of kritsuk11. The analyses and simulations presented in this paper were performed on the NASA Pleiades cluster supported by the NASA High End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center and on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University. Support for V.S. was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51445.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555, and by Harvard University through the Institute for Theory and Computation Fellowship. The analyses presented in this paper were greatly aided by the following free software packages: NumPy (numpy_ndarray), SciPy (scipy), Matplotlib (matplotlib), and yt (yt). We have also used the Astrophysics Data Service (ADS) and arXiv preprint repository extensively during this project and the writing of the paper.
@@ Appendix A Dependence on Resolution
The turbulent cascade in simulations without explicit viscosity truncates at the numerical dissipation scale close to the resolution. As Figure 7 shows, in our high-resolution ($1024^{3}$) DNS that we use to calibrate and test our SLES model, this truncation starts around $k/k_{\rm min}=L_{\rm box}/L\sim 100$, which corresponds to $\sim$10 cells, in agreement with the kritsuk11 results. Only the estimates of small-scale turbulence, $e_{\rm turb}$, on sufficiently larger scales should be used for comparisons with SLES, as $e_{\rm turb}$ on smaller scales is not numerically converged.
In Figure 12, we investigate the appropriate coarse-graining scales that can be used for our purposes. Different line styles show DNSs with different resolutions, while the line colors show the average $E_{\rm turb}\equiv\langle e_{\rm turb}\rangle_{\rm box}$ coarse-grained using different coarse grids: $16^{3}$, $32^{3}$, $64^{3}$, and $128^{3}$ (see Section 3.1). A finer coarse grid corresponds to the estimate of $e_{\rm turb}$ over a smaller number of mesh cells, e.g., for our fiducial $1024^{3}$ DNS, these grids correspond to 64, 32, 16, and 8 cells per side.
The comparison of DNSs with different resolutions and spacing between results for different coarse grids shows that the results of our DNS run are converged when considered on down to $64^{3}$ coarse grids. In Section 3, however, we also show the results for $128^{3}$, which should be interpreted with caution as they probe $e_{\rm turb}$ on scales close to the numerical dissipation scale.